OpenStudy (alex_mattucci):

Please Help Me! Transform x2 + 10x + 24 = 0 into the form (x - p)2 = q? [p and q are integers] Three steps only! Show Work!

OpenStudy (alex_mattucci):

Getting answer in a sec

OpenStudy (anonymous):

did u get it answer

OpenStudy (alex_mattucci):

no sorry, it won't get them posted

OpenStudy (anonymous):

ok

OpenStudy (alex_mattucci):

can you walk me through a step by step of this

OpenStudy (alex_mattucci):

preferably 3 steps

OpenStudy (anonymous):

the last one for sure 1st not can't be cuz 0x + 24 = 0 can't positively = out (x - p)2 = q ID think so

OpenStudy (alex_mattucci):

What?

OpenStudy (anonymous):

just nvm

OpenStudy (anonymous):

@Daniellelovee por favor ayudenos

OpenStudy (anonymous):

does it make sense no

OpenStudy (alex_mattucci):

no

OpenStudy (alex_mattucci):

@Directrix Please help here

zepdrix (zepdrix):

So this problem is asking you to `complete the square`.\[\large\rm x^2+10x+24=0\]Subtracting 24,\[\large\rm x^2+10x\qquad\qquad=-24\]We'll attempt to complete the square on these x's.

zepdrix (zepdrix):

When we have a quadratic in this form: \(\large\rm x^2+bx\) There is a special value we want to add to the expression to complete it. That value is calculated this way: You take your b coefficient, cut it half, then square it.\[\large\rm \left(\frac{b}{2}\right)^2\]

zepdrix (zepdrix):

So for this problem, we see that 10 is the b coefficient,\[\large\rm \left(\frac{10}{2}\right)^2=25\]Cut 10 in half, and then squaring it gives us 25. This is the special value that we'll add to both sides of our equation to complete the square.

zepdrix (zepdrix):

\[\large\rm x^2+10x+25=-24+25\]

zepdrix (zepdrix):

\[\large\rm x^2+10x+25=1\]

zepdrix (zepdrix):

Then our expression on the left will simplify down to \(\large\rm \left(x+\frac{b}{2}\right)^2\) You can verify this by doing factoring by grouping from this point. But anyway, should get something like this,\[\large\rm (x+5)^2=1\]

OpenStudy (alex_mattucci):

ok, i got that so far!

OpenStudy (alex_mattucci):

oh, that is the answer!

zepdrix (zepdrix):

From this point, compare this to the expression involving p and q, and try to match up the numbers carefully.

zepdrix (zepdrix):

Or did they just want it in that form :) Ya maybe I misread the question hehe

OpenStudy (alex_mattucci):

Thank you so much for your help, could you possibly help me with one more problem though?